toms weight plus twice my weight is 361. My weight plus twice toms weight is 362. How much is my weight and toms weight?
Answer by checkley77(12844)

t+2m=361
t=361-2m
2t+m=362
Substitute 361-2m for t in the second equation.
2(361-2m)+m=362
722-4m+m=362
-3m=362-722
-3m=-360
m=-360/-3
m=120 lbs. for my weight.
t+2*120=361
t+240=361
t=361-240
t=121 lbs for Tom's weight.